draw a circle of r 4cm.from a point p,7 from the centre of the circle,draw a pt of tangents to the circle measure length of each tangent

Grade:8

2 Answers

unknown
7 years ago
  1. draw a circle of radii 4cm and name the centre as o
  2. take a point p as according to question
  3. draw perpendicular bisector of op. and name the point as q.
  4. taking q as centre,draw a circle taking radii qp=oq
  5. the point at which this circle touch the older one name it as A and B
JOIN AP and BP,which are the required tangents
 hope u understood well...........
Saravanan
7 years ago
i can give justification
 PA & PB are tangents from p to circle  from p 
                OP is dia.m
              angle oap is 9o degeree
                OA perpendicular to AP
 we know that,radius is perpendicular to tangent 
            PA is tangent 
         similarly,AB is tangent
        in traingle OAP,
     angle A=90 degree (by.py.th)
       PA=OP-OA the whole sq
           =49-16=root of 33cm